# Axis-Aligned Document Dewarping

### v1

@@ Abstract
Document dewarping is crucial for many applications. However, existing learning–based methods primarily rely on supervised regression with annotated data without leveraging the inherent geometric properties in physical documents to the dewarping process. Our key insight is that a well-dewarped document is characterized by transforming distorted feature lines into axis‐aligned ones. This property aligns with the inherent axis‐aligned nature of the discrete grid geometry in planar documents. In the training phase, we propose an axis-aligned geometric constraint to enhance document dewarping. In the inference phase, we propose an axis alignment preprocessing strategy to reduce the dewarping difficulty. In the evaluation phase, we introduce a new metric, Axis-Aligned Distortion (AAD), that not only incorporates geometric meaning and aligns with human visual perception but also demonstrates greater robustness. As a result, our method achieves SOTA results on multiple existing benchmarks and achieves $\boldsymbol{\sim}$34.5improvements on the AAD metric.
@@ 1 Introduction
The digitization of documents dramatically facilitates our lives by converting information on physical documents into electronic data that can be used with digital devices. However, due to paper deformation, scanners must be fixed in position for precise scanning and acquisition, so this method has specific limitations. With the widespread use of mobile phones and cameras, photography has become a primary tool for digitization. Nonetheless, images taken by handheld cameras often suffer from geometric distortions and uneven lighting, negatively impacting the user’s reading experience and downstream tasks like optical character recognition (OCR).
@@
In early research, many studies defined this problem as surface reconstruction from different imaging devices, including multi-view images [33, 19], stereo cameras [34], and structured light cameras [1, 26]. After that, some studies have leveraged document priors in single-view images to conduct extensive research, such as document boundaries [2], text lines [13, 14]. However, these are affected by the low-level features of the detected prior knowledge.
@@
In the years of the flourishing development of deep learning, convolutional neural networks began to learn 2D deformation field mapping from deformed images to expected dewarping results [23, 5, 39, 7, 24, 17, 16, 8, 12, 41]. Data-driven methods perform supervised learning dewarping under the guidance of various supervision signals with synthetic annotations [23, 36, 17], such as depth [5], geometric semantics [25], control points [40, 36, 41], layout [17].
@@
Although prior methods perform well on simple deformations, they often fail on document images with complex deformations. This is mainly because previous studies rely on strong supervision signals, such as control points, which are effective but have no geometric meaning, and text lines, which are meaningful but difficult to extract and lack generalization. These studies do not effectively use the inherent geometric information in the paper documents, which is investigated and exploited in this paper.
Figure 1: The top row is the warped local document image, and the bottom is the expected dewarping result. The main change feature is that the feature line aligns with the axis as much as possible. The dotted lines are added for better visibility.
In this paper, we perform document dewarping using a lightweight convolutional neural network that predicts a 2D unwarping grid from the distorted image. On this basis, to make the network better learn the nature of document dewarping, as shown in the dewarping example in Fig. 1, we expect the feature lines in the dewarped document to be more horizontally or vertically aligned. We leverage the inherent axis alignment property of the discrete grid geometry in planar documents, design a corresponding axis-aligned geometric constraint, and integrate it into the network training process. This dramatically improves the dewarping performance of the network. In the inference stage, to improve the dewarping performance for images containing target documents in different proportions and rotation angles, we propose an axis alignment preprocessing strategy. This strategy simplifies the dewarping process by leveraging the position properties of grid points and the axis alignment properties of planar document geometry. In the evaluation stage, we propose a new evaluation metric, AAD, to automatically measure the axis alignment distortion of the dewarped results, which have specific geometric meaning and align with human visual perception.
@@ 2 Related Work
Current research on document dewarping can be broadly categorized into traditional optimization methods and deep learning-based methods. As mentioned in the Sec. 1, traditional approaches to document dewarping primarily involve image processing and 3D reconstruction, preceding the advent of deep learning.
The research on image processing is to model the most commonly used low-level features extracted from document images, including letters/words, boundaries, spacing, and text lines, and to correct perspective and page curl distortion through local line spacing and cell shape, which can be formulated as an energy minimization optimization problem [35, 30, 14, 31]. In the recent research, Jiang et al. [12] applied deep learning to extract boundaries and document lines, followed by geometric constraint optimization. However, it is not easy to detect and perceive low-level features in deformed document images, which limits the application of methods based on image processing.
In 3D reconstruction research, variations in shadows from a single image or across different imaging devices, along with other input cues, have been utilized to reconstruct 3D surfaces [32, 37, 43, 19, 9]. Some studies model these surfaces as special types, such as cylindrical surfaces [3] and developable surfaces [26]. Recently, Luo et al. [22] treat the document as a developable surface and employ a recently proposed isometric constraint for modeling [11]. However, these methods rely on detecting document feature lines, resulting in long optimization times and limited generality.
In the research of deep learning, the current research has generated a large amount of data using rendering techniques and dataset collection, such as Doc3D [5], Inv3D [10], Simulated paper [17], UVDoc [36]. These data-driven approaches can be categorized into forward mapping methods [23, 18, 12] and backward mapping methods [5, 6, 7, 8]. To further improve optimization, some methods incorporate additional supervision signals, such as text lines, masks, boundaries, layouts, angles, and curvatures to further improve optimization [42, 7, 4, 16, 17]. However, extracting these additional cues can introduce extra noise and lack generalizability.
An effective approach is to fully exploit the inherent geometric properties of the document, which are independent of annotated labels and generally more robust. As Xie et al. [39] exploited the continuity property of document deformation by incorporating local fairing constraints into the displacement prediction process. Furthermore, the recent GSO-Net [38], based on grid surface modeling, achieves accurate and robust optimization by learning the intrinsic geometric properties.
@@ 3 Method
We adopt the convolutional neural network architecture from UVDoc [36] as the foundation for our document dewarping network. This fully convolutional deep neural network simultaneously predicts the document’s 3D grid mesh and corresponding 2D unwarping grid in a dual-task framework, which is inspired by the work of Wang et al. [39, 40]
In the following sections, Sec. 3.1 details the axis-aligned geometric constraint introduced during training, Sec. 3.2 analyzes the loss functions employed during network training, and Sec. 3.3 describes the preprocessing strategy for axis alignment during inference.
@@ 3.1 Axis-Aligned Geometric Constraints
Figure 2: Axis-aligned geometric constraints were enforced during document dewarping network training. The top row shows the predicted 2D unwarping grid (obtained from Grid-Net in UVDoc [36]), and the bottom row displays its transformation into UV space based on the ground truth, facilitating the computation of the corresponding axis-aligned geometric constraints loss along horizontal and vertical directions.
Fig. 2 illustrates the pipeline for computing the axis-aligned geometric constraint loss during training. Our approach takes advantage of the inherent geometric properties of planar paper, where the corresponding discrete grid geometry is aligned along the horizontal and vertical axes (as shown by $Q^{GT}$ in Fig. 2). This property aligns with the horizontal and vertical orientations of the document’s feature lines in image space. So, we can assess the axis alignment distortion of these feature lines in the dewarped image by computing the corresponding deviation of the unwrapped grid in planar space. However, since the current predictions are in image space, directly calculating this error is challenging.
To address this issue, we first employ an interpolation function to map the predicted 2D unwarping grid from image space $P$ to UV space $Q$. In the UV space, the deviation from horizontal and vertical alignment is effectively quantified by computing the position variance along each row and column. By integrating these axis-aligned geometric constraints into the training process, we can guide the network in learning to dewarp distorted document feature lines toward the axis-aligned direction. The following are the details.
@@
During training, the corresponding ground truth grid is,:
@@
$q=f(p)=\mathrm{Interpolate}\bigl{(}q^{GT};\,p^{GT},\,P\bigr{)}$ (4)
@@
$\mathcal{L}_{\text{ver}}=\sum\nolimits_{i=1}^{w}\mathrm{Var}\Big{(}\{u_{i,1},u_{i,2},\dots,u_{i,h}\}\Big{)}.$ (7)
@@
Minimizing $\mathcal{L}_{\text{AL}}$ means the predicted 2D unwarping grid after being mapped into the UV space, as much as possible to align with the axis. Thus, the feature lines on distorted image dewarping results are axis-aligned as much as possible.
@@ 3.2 Loss Function
Grid 2D and 3D Losses. Following UVDoc [36], we supervise the network with the losses:
$\mathcal{L}_{2D}=\|G_{2D}-G_{2D}^{\text{gt}}\|_{1},\mathcal{L}_{3D}=\|G_{3D}-G_{3D}^{\text{gt}}\|_{1}$ (9)
@@
Axis-Aligned Geometric Constraint Loss. We impose axis-aligned geometric constraints $\mathcal{L}_{AL}$ in the UV space, guiding the network to achieve effective axis alignment dewarping of feature lines, as depicted in Sec. 3.1.
@@ 3.3 Axis Alignment Preprocessing for Inference
During inference, previous methods employ additional segmentation and deformation models for preprocessing [24, 7] to improve dewarping robustness under diverse input conditions. Essentially, these methods distinguish between the background and the document region, enabling the model to focus on the document area and reducing dewarping difficulty. However, they do not directly address the elimination of geometric distortion.
Drawing on the insights from Fig. 1, axis alignment is central to the dewarping process. So, we propose an axis alignment preprocessing strategy, as illustrated in Fig. 3. For a given input image, we first compute the minimum‐area rotation rectangle using the positional information from the predicted 2D unwarping grid. Subsequently, we perform an axis alignment rotation and clip the target area to obtain a preprocessed result. We then reapply the inference process on this preprocessed image to yield a refined dewarping outcome. If necessary, this process can be iteratively repeated to achieve further improvement under varying conditions. As shown in Fig. 3, the fine results obtained after preprocessing outperform the coarse results, effectively reducing the influence of the target position on the overall correction.
@@
In contrast to previous preprocessing methods [24, 7], our approach eliminates the need for additional segmentation or deformation models. Instead, it directly leverages the predicted 2D unwarping grid to extract object region information, rendering the method more efficient and streamlined. A detailed analysis is provided in Sec. 4.5.
@@ 4.1 Datasets and Benchmarks
4.1.1 Training Datasets
We used two datasets, Doc3D [5] and UVDoc [36], for alternative optimization models following the setting of Verhoeven et al. [36].
Doc3D [5] is the largest distortion dewarping dataset, containing 100K distorted images. It was created using approximately 4,000 authentic document images and rendering software. The labels for each distorted document image include 3D coordinate maps, albedo maps, normals, depth maps, UV maps, and backward mapping maps.
UVDoc [36] aims to reduce the domain gap between the synthetic Doc3D dataset [5]—commonly used for training document unwarping models—and real document photographs. This dataset contains 20K pseudo-photorealistic document images and provides all the necessary information to train a coarse grid-based document unwarping neural network.
4.1.2 Benchmarks
We evaluated our method on the following three benchmarks:
DocUNet [23]: This benchmark comprises 130 distorted images captured in natural scenes using mobile devices. Following previous methods, 25 rich-text images were selected for OCR evaluation, and according to DocGeoNet [8], the 127th and 128th images (which did not match the labels) were rotated by 180.
DIR300 [8]: Captured with a moving camera, this benchmark contains 300 images of real distorted documents with complex backgrounds, varying degrees of distortion, and diverse lighting conditions. In line with previous studies, 90 rich-text images were selected for OCR evaluation.
UVDoc [36]: Generated via a pseudo-photorealistic rendering pipeline, this benchmark consists of 50 images, including matching pairs with and without illumination. Following the UVDoc protocol, 20 rich-text images were selected for OCR evaluation.
@@ 4.2 Evaluation Metrics
MS-SSIM, LD and AD. Following Ma et al. [23], we employ multi-scale structural similarity (MS-SSIM) and local distortion (LD) to assess image similarity. Additionally, we use the aligned distortion (AD) metric introduced in [24] to overcome the limitations of earlier measures.
CER and ED. Consistent with previous work, we measure OCR performance using Character Error Rate (CER) [27] and Edit Distance (ED) [15]. Our experiments utilize OCR recognition with Tesseract v5.4.0.20240606 and PyTesseract v0.3.13  [^1] https://pypi.org/project/pytesseract/.
@@
4.2.1 AAD Metric
As illustrated in Fig. 1, our visual evaluation of document dewarped primarily relies on the axis alignment error of the document’s feature lines, which carries clear geometric significance. However, existing metrics fail to capture this characteristic effectively. We introduce the AAD (Axis-Aligned Distortion) metric, which quantifies distortion by measuring the axial alignment error of the dewarped results. The metric meaning is the same as in 3.1, the key difference is that this metric is computed during the evaluation phase using given image pairs rather than 2D grid pairs from the training phase.
Let $y$ denote the ground truth image. The optical flow field from $y$ to its dewarped version is computed using the SIFT-flow algorithm [20], yielding flow components $v_{x}$ and $v_{y}$ along the horizontal and vertical directions, respectively. The directional gradients of $y$ are then extracted using the Sobel operator [29] along each axis and normalized:
@@ 4.2.1 AAD Metric
Here, $m_{i}$ and $n_{j}$ denote the gradient-weighted mean of the vertical flow in the $i$th row and the horizontal flow in the $j$th column, respectively, while $d^{\text{row}}_{i,j}$ and $d^{\text{col}}_{i,j}$ measure the corresponding deviations in pixel ($i\,j$).
@@
Fig. 4 shows an example visualization of the AAD metric. For two different dewarped results, the alignment of table feature lines and the horizontal alignment of text lines serve as the primary evaluation criteria. By overlaying the horizontal and vertical components of the AAD heatmap onto the dewarped results, the corresponding axis alignment errors are displayed in detail. The heatmap colors correlate well with the calculated values, offering a perceptual measure of axis alignment.
@@
Figure 4: Example of visualizing AAD metrics. The first row is the ground truth image and its gradient heatmap in the horizontal ($G_{y}$) and vertical ($G_{x}$) directions; the second and third rows correspond to the two different dewarped results and their overlay heatmaps for AAD and its horizontal (AAD_H) and vertical (AAD_V) components, it can according to the $\updownarrow$ and $\leftrightarrow$ arrows to see the distortion error in the alignment direction of the corresponding axes.
To validate the robustness of the proposed AAD metric, we applied geometric and illumination disturbances to the dewarped images shown in Fig. 5, generating 100 warped images for each setting. The known disturbance serves as the optical flow information for calculating the ground truth metrics, while the optical flow information obtained via the SIFT-flow[20] method is used to compute the reference metrics. The linear correlation between the reference value and the actual value of each metric can be a criterion for their robustness. As shown in Fig. 6(a), the AAD metric demonstrated the highest linear correlation ($R^{\text{2}}$) across all sets, indicating a strong linear relationship with the ground truth even under minimal geometric disturbances(i.e., in regions with lower metric values, which are critical for comparing dewarped results that closely approximate the true values), In contrast, the AD metric exhibited a nonlinear response in these regions. In Fig. 6(b), we evaluated the normalized standard deviation under three different illumination conditions for each image in Fig. 5. The results revealed that the standard deviation for AAD (0.0230) was smaller than that for AD (0.0252), further confirming its enhanced robustness.
Figure 5: Geometric and illumination disturbances applied to the dewarped image. On the left is the distorted corresponding to the dewarped and ground truth images. On the right, geometric and illumination disturbances are added to the dewarped image, and the intensity increases along with the corresponding arrow direction. Set1 includes only geometric disturbances, Set2 adds color interference to Set1, and Set3 adds a random shadow to Set2.
Figure 6: Robustness analysis of each metric. (a) The x-axis represents the ground truth metric computed from the ground truth optical flow, while the y-axis represents the reference metric derived from the SIFT-flow method; ideally, they should exhibit a high linear correlation. (b) The normalized standard deviation of the metrics under three different environmental conditions, lower values indicate greater robustness to environmental disturbances.

### v2

@@ Abstract
Document dewarping is crucial for many applications. However, existing learning-based methods rely heavily on supervised regression with annotated data without fully leveraging the inherent geometric properties of physical documents. Our key insight is that a well-dewarped document is defined by its axis-aligned feature lines. This property aligns with the inherent axis-aligned nature of the discrete grid geometry in planar documents. Harnessing this property, we introduce three synergistic contributions: for the training phase, we propose an axis-aligned geometric constraint to enhance document dewarping; for the inference phase, we propose an axis alignment preprocessing strategy to reduce the dewarping difficulty; and for the evaluation phase, we introduce a new metric, Axis-Aligned Distortion (AAD), that not only incorporates geometric meaning and aligns with human visual perception but also demonstrates greater robustness. As a result, our method achieves state-of-the-art performance on multiple existing benchmarks, improving the AAD metric by 18.2% to 34.5%. The code is publicly available at https://github.com/chaoyunwang/AADD.
@@ 1 Introduction
The digitization of documents dramatically facilitates our lives by converting information on physical documents into electronic data that can be used with digital devices. However, traditional scanners, which require fixed positioning for precise acquisition, have limitations when dealing with paper deformation. With the widespread use of mobile phones and cameras, photography has become a primary tool for digitization. Nonetheless, images captured by handheld cameras often suffer from geometric distortions and uneven lighting, which degrade the reading experience and hinder downstream tasks like optical character recognition (OCR).
@@
In early research, many studies defined this problem as surface reconstruction from different imaging devices, including multi-view images (Tsoi and Brown 2007; Liang, DeMenthon, and Doermann 2008), stereo cameras (Ulges, Lampert, and Breuel 2004), and structured light cameras (Brown and Seales 2001; Meng et al. 2014). Subsequent studies on single-view images have leveraged various document priors, such as document boundaries (Brown and Tsoi 2006), text lines (Kil et al. 2017; Kim, Koo, and Cho 2015). However, the performance of these methods is often limited by the robustness of low-level feature detection.
@@
With the recent development of deep learning, convolutional neural networks began to learn 2D deformation field mapping from deformed images to expected dewarping results (Ma et al. 2018; Das et al. 2019; Xie et al. 2020; Feng et al. 2021; Ma et al. 2022; Li et al. 2023b, a; Feng et al. 2022; Jiang et al. 2022; Yu et al. 2024). Data-driven methods perform supervised learning dewarping under the guidance of various supervision signals with synthetic annotations (Ma et al. 2018; Li et al. 2023b; Verhoeven, Magne, and Sorkine-Hornung 2023), such as depth (Das et al. 2019), geometric semantics (Markovitz et al. 2020), control points (Verhoeven, Magne, and Sorkine-Hornung 2023; Xie et al. 2021; Yu et al. 2024), layout (Li et al. 2023b).
@@
Although prior methods perform well on simple deformations, they often fail on document images with complex deformations. This is mainly because previous studies rely on strong supervision signals, such as control points, which are effective but have no geometric meaning, and text lines, which are meaningful but difficult to extract and often exhibit poor generalization. These methods often overlook the fundamental geometric properties of paper documents—a gap our work aims to fill.
Figure 1: Our research motivation and main contributions. (a) A warped document (left) and its target rectified version (right). The key characteristic of the rectified version is the alignment of its feature lines (highlighted by dotted lines) with the axes. We term this the “axis-aligned property”. (b) Inspired by (a), we integrate this axis-aligned property into the training, inference, and evaluation stages of our deep learning method.
In this paper, we perform document dewarping using a lightweight convolutional neural network that predicts a 2D unwarping grid from the distorted image. Our key insight, as illustrated in Figure 1(a), is that the defining characteristic of a successfully rectified document is the alignment of its feature lines with the coordinate axes. We term this fundamental observation the “axis-aligned property”—a principle rooted in the inherent discrete grid geometry of planar documents.
Inspired by this single, powerful principle, we systematically integrate it into all three key stages of our deep learning pipeline, as outlined in Figure 1(b): For training, we design an axis-aligned geometric constraint that explicitly guides the network to learn this property, significantly improving dewarping performance. For inference, we introduce a novel axis-alignment preprocessing strategy. This strategy leverages the property to normalize the document’s orientation beforehand, simplifying the task for the network and boosting its robustness against various rotations and scales. Finally, for evaluation, we propose a new metric, Axis-Aligned Distortion (AAD), which directly quantifies how well the dewarped result adheres to the axis-aligned property, offering a measure that is both geometrically meaningful and consistent with human perception.
@@ 2 Related Work
Current research on document dewarping can be broadly categorized into traditional optimization methods and deep learning-based methods. As mentioned in Section 1, traditional approaches to document dewarping primarily involve image processing and 3D reconstruction, preceding the advent of deep learning.
Research in image processing has focused on modeling the most commonly used low-level features extracted from document images, including letters/words, boundaries, spacing, and text lines, and to correct perspective and page curl distortion through local line spacing and cell shape, which can be formulated as an energy minimization optimization problem (Ulges, Lampert, and Breuel 2005; Stamatopoulos et al. 2010; Kim, Koo, and Cho 2015; Takezawa, Hasegawa, and Tabbone 2017). More recently, Jiang et al. (Jiang et al. 2022) applied deep learning to extract boundaries and document lines, followed by geometric constraint optimization. However, reliably detecting such low-level features in heavily distorted images is challenging, limiting the applicability of these methods.
In 3D reconstruction research, variations in shadows from a single image or across different imaging devices, along with other input cues, have been utilized to reconstruct 3D surfaces (Tan et al. 2005; Wada, Ukida, and Matsuyama 1997; Zhang et al. 2009; Liang, DeMenthon, and Doermann 2008; He et al. 2013). Some studies model these surfaces as special types, such as cylindrical surfaces (Cao, Ding, and Liu 2003) and developable surfaces (Meng et al. 2014). Recently, Luo et al. (Luo and Bo 2022) treat the document as a developable surface and employ a recently proposed isometric constraint for modeling (Jiang et al. 2020). However, these methods rely on detecting document feature lines, resulting in long optimization times and limited generality.
Deep learning methods leverage large amounts of data generated by rendering techniques and dataset collection, such as Doc3D (Das et al. 2019), Inv3D (Hertlein, Naumann, and Philipp 2023), Simulated paper (Li et al. 2023b), UVDoc (Verhoeven et al. 2023). These data-driven approaches can be categorized into forward mapping methods (Ma et al. 2018; Li et al. 2019; Jiang et al. 2022) and backward mapping methods (Das et al. 2019, 2021; Feng et al. 2021, 2022). To further improve optimization, some methods incorporate additional supervision signals, such as text lines, masks, boundaries, layouts, angles, and curvatures to further improve optimization (Zhang et al. 2022; Feng et al. 2021; Dai et al. 2023; Li et al. 2023a, b). However, extracting these additional cues can introduce extra noise and lack generalizability.
An effective approach is to fully exploit the inherent geometric properties of the document, which are independent of annotated labels and generally more robust. As Xie et al. (Xie et al. 2020) exploited the continuity property of document deformation by incorporating local fairing constraints into the displacement prediction process. Furthermore, the recent work (Wang et al. 2024, 2025), based on grid surface modeling, achieves accurate and robust optimization by learning the intrinsic geometric properties.
@@ 3 Method
We adopt the convolutional neural network architecture from UVDoc (Verhoeven et al. 2023) as the foundation for our document dewarping network. This fully convolutional deep neural network simultaneously predicts the document’s 3D grid mesh and corresponding 2D unwarping grid in a dual-task framework, which is inspired by the work of Xie et al. (Xie et al. 2020, 2021)
In the following sections, Section 3.1 details the axis-aligned geometric constraint introduced during training, Section 3.2 analyzes the loss functions employed during network training, and Section 3.3 describes the preprocessing strategy for axis alignment during inference.
@@ 3.1 Axis-Aligned Geometric Constraints
Figure 2: Axis-aligned geometric constraints were enforced during document dewarping network training. The top row shows the predicted 2D unwarping grid (obtained from Grid-Net in UVDoc (Verhoeven et al. 2023)), and the bottom row displays its transformation into UV space based on the ground truth, facilitating the computation of the corresponding axis-aligned geometric constraints loss along horizontal and vertical directions.
Figure 2 illustrates the pipeline for computing the axis-aligned geometric constraint loss during training. Our approach takes advantage of the inherent geometric properties of planar paper, where the corresponding discrete grid geometry is aligned along the horizontal and vertical axes (as shown by $Q^{GT}$ in Figure 2). This property aligns with the horizontal and vertical orientations of the document’s feature lines in image space. Therefore, we can assess the axis alignment distortion of these feature lines in the dewarped image by computing the corresponding deviation of the unwrapped grid in planar space. However, since the current predictions are in image space, directly calculating this error is challenging.
To address this issue, we first employ an interpolation function to map the predicted 2D unwarping grid from image space $P$ to UV space $Q$. In the UV space, the deviation from horizontal and vertical alignment is effectively quantified by computing the position variance along each row and column. By integrating these axis-aligned geometric constraints into the training process, we can guide the network in learning to dewarp distorted document feature lines toward the axis-aligned direction.
@@
During training, the corresponding ground truth grid is:
@@
$q=f(p)=\mathrm{Interpolate}\bigl(q^{GT};\,p^{GT},\,P\bigr).$ (4)
@@
$\mathcal{L}_{\text{ver}}$ $=\sum\nolimits_{i=1}^{w}\mathrm{Var}\Big(\{u_{i,1},u_{i,2},\dots,u_{i,h}\}\Big).$
@@
Minimizing $\mathcal{L}_{\text{AL}}$ encourages the predicted 2D unwarping grid, when mapped to the UV space, to align with the axes as closely as possible.
@@ 3.2 Loss Function
Grid 2D and 3D Losses. Following UVDoc (Verhoeven et al. 2023), we supervise the network with the losses:
$\mathcal{L}_{2D}=\|G_{2D}-G_{2D}^{\text{gt}}\|_{1},\mathcal{L}_{3D}=\|G_{3D}-G_{3D}^{\text{gt}}\|_{1}.$ (8)
@@
Axis-Aligned Geometric Constraint Loss. We impose axis-aligned geometric constraints $\mathcal{L}_{AL}$ in the UV space, guiding the network to achieve effective axis alignment dewarping of feature lines, as depicted in Section 3.1.
@@ 3.3 Axis Alignment Preprocessing for Inference
During inference, previous methods employ additional segmentation and deformation models for preprocessing (Ma et al. 2022; Feng et al. 2021) to improve dewarping robustness under diverse input conditions. Essentially, these methods distinguish between the background and the document region, enabling the model to focus on the document area and reducing dewarping difficulty. However, they do not directly address the elimination of geometric distortion.
Drawing on the insights from Figure 1(a), axis alignment is central to the dewarping process. So, we propose an axis alignment preprocessing strategy, as illustrated in Figure 3. For a given input image, we first compute the minimum‐area rotation rectangle using the positional information from the predicted 2D unwarping grid. Subsequently, we rotate the image to align the document’s principal axes and clip the target area. This preprocessed image is then fed back into the network for a refined prediction. If necessary, this process can be iteratively repeated to achieve further improvement under varying conditions. As shown in Figure 3, the fine results obtained after preprocessing outperform the coarse results, effectively reducing the influence of the target position on the overall correction.
@@
Unlike prior methods that rely on external segmentation models (Ma et al. 2022; Feng et al. 2021), our preprocessing is self-contained. It directly leverages the predicted 2D unwarping grid to extract object region information, rendering the method more efficient and elegant. A detailed analysis is provided in Section 4.5.
@@ 4.1 Datasets and Benchmarks
Training Datasets
We used two datasets, Doc3D (Das et al. 2019) and UVDoc (Verhoeven et al. 2023), for alternative optimization models following the setting of (Verhoeven et al. 2023).
Doc3D (Das et al. 2019) is the largest distortion dewarping dataset, containing 100K distorted images. It was created using approximately 4,000 authentic document images and rendering software. The labels for each distorted document image include 3D coordinate maps, albedo maps, normals, depth maps, UV maps, and backward mapping maps.
UVDoc (Verhoeven et al. 2023) aims to reduce the domain gap between the synthetic Doc3D dataset (Das et al. 2019)—commonly used for training document unwarping models—and real document photographs. This dataset contains 20K pseudo-photorealistic document images and provides all the necessary information to train a coarse grid-based document unwarping neural network.
Benchmarks
We evaluated our method on the following two benchmarks:
DocUNet (Ma et al. 2018): This benchmark comprises 130 distorted images captured in natural scenes using mobile devices. Following previous methods, 25 rich-text images were selected for OCR evaluation, and according to DocGeoNet (Feng et al. 2022), the 127th and 128th images (which did not match the labels) were rotated by 180.
DIR300 (Feng et al. 2022): Captured with a moving camera, this benchmark contains 300 images of real distorted documents with complex backgrounds, varying degrees of distortion, and diverse lighting conditions. In line with previous studies, 90 rich-text images were selected for OCR evaluation.
@@ 4.2 Evaluation Metrics
MS-SSIM, LD and AD. Following Ma et al. (Ma et al. 2018), we employ multi-scale structural similarity (MS-SSIM) and local distortion (LD) to assess image similarity. Additionally, we use the aligned distortion (AD) metric introduced in (Ma et al. 2022) to overcome the limitations of earlier measures.
CER and ED. Consistent with previous work, we measure OCR performance using Character Error Rate (CER) (Morris, Maier, and Green 2004) and Edit Distance (ED) (Levenshtein 1966). Our experiments utilize OCR recognition with Tesseract v5.4.0.20240606 and PyTesseract v0.3.13  [^1] https://pypi.org/project/pytesseract/.
@@
AAD Metric
As illustrated in Figure 1(a), our visual evaluation of dewarped documents primarily relies on the axis alignment error of the document’s feature lines, which carries clear geometric significance. However, existing metrics fail to capture this characteristic effectively. We introduce the AAD (Axis-Aligned Distortion) metric, which quantifies distortion by measuring the axial alignment error of the dewarped results. It is conceptually identical to our geometric constraint (Section 3.1), with the key difference being its application: it is computed on image pairs during evaluation, rather than on 2D grid pairs during training.
Let $y$ denote the ground truth image. The optical flow field from $y$ to its dewarped version is computed using the SIFT-flow algorithm (Liu, Yuen, and Torralba 2010), yielding flow components $v_{x}$ and $v_{y}$ along the horizontal and vertical directions, respectively. The directional gradients of $y$ are then extracted using the Sobel operator (Sobel, Feldman et al. 1968) along each axis and normalized:
@@
Here, $m_{i}$ and $n_{j}$ denote the gradient-weighted mean of the vertical flow in the $i$th row and the horizontal flow in the $j$th column, respectively, while $d^{\text{row}}_{i,j}$ and $d^{\text{col}}_{i,j}$ measure the corresponding deviations in pixel ($i,j$).
@@
Figure 4: Example of visualizing AAD metrics. The first row is the ground truth image and its gradient heatmap in the horizontal ($G_{y}$) and vertical ($G_{x}$) directions; the second and third rows correspond to the two different dewarped results and their overlay heatmaps for AAD and its horizontal (AAD_H) and vertical (AAD_V) components, the $\leftrightarrow$ and $\updownarrow$ arrows indicate the directions for visualizing the axis-aligned error of the dewarped feature lines.
Figure 4 shows an example visualization of the AAD metric. For two different dewarped results, the alignment of table feature lines and the horizontal alignment of text lines serve as the primary evaluation criteria. By overlaying the horizontal (row) and vertical (col) components of the AAD heatmap onto the dewarped results, the corresponding axis alignment errors are displayed in detail. The heatmap colors correlate well with the calculated values, offering a perceptual measure of axis alignment.
Figure 5 illustrates the superiority of our proposed AAD metric over the existing AD metric. In (b), our dewarped result is visibly superior to that of UVDoc (Verhoeven et al. 2023) due to better text-line alignment, although the two appear broadly similar. The AD metric, however, fails on two counts: its heatmap in (c) is uninterpretable, and its numerical score contradicts this visual reality by rating our result as worse. In contrast, our AAD metric is founded on a clear geometric principle. This gives its heatmap in (d) a highly interpretable meaning, where brighter areas directly correspond to distorted feature lines. As a result, both its visual heatmap and numerical score correctly identify our result as superior, which is consistent with human judgment. This ability to distinguish subtle differences is essential for guiding the next stage of research as performance gaps between state-of-the-art models narrow.
Table 1: Quantitative evaluation and ablation study on the DocUNet and DIR300 benchmarks. We compare our method with previous methods and ablate our components: Axis-aligned geometric constraint Loss (AL) and Axis alignment Preprocessing (AP). “↑” indicates higher is better, “↓” means lower is better. Bold denotes the best result, and underline indicates second-best. The “Improvements” row shows the relative gain of our full model over the previous state-of-the-art.
Method DocUNet benchmark DIR300 benchmark
MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓ MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓
DewarpNet(Das et al. 2019) 0.474 8.362 0.398 0.164 824.5 0.225 0.492 13.944 0.332 0.147 1076.8 0.336
DDCP(Xie et al. 2021) 0.472 8.982 0.428 0.159 1124.7 0.280 0.552 11.02 0.357 0.132 739.3 0.257
DocTr(Feng et al. 2021) 0.509 7.773 0.369 0.151 708.6 0.185 0.616 7.189 0.255 0.107 698.4 0.211
PaperEdge(Ma et al. 2022) 0.473 7.967 0.368 0.121 731.4 0.183 0.584 7.968 0.255 0.091 447.4 0.177
DocGeoNet(Feng et al. 2022) 0.504 7.717 0.381 0.158 695.0 0.187 0.638 6.403 0.241 0.106 672.5 0.206
FTDR(Li et al. 2023a) 0.497 8.416 0.377 0.151 688.3 0.176 0.607 7.680 0.243 0.108 639.4 0.198
LADoc(Li et al. 2023b) 0.525 6.706 0.300 0.121 689.8 0.180 0.652 5.702 0.195 0.087 495.4 0.173
UVDoc(Verhoeven et al. 2023) 0.545 6.827 0.316 0.125 754.2 0.193 0.621 7.730 0.219 0.101 614.0 0.237
Ours (Baseline, w/o AL, AP) 0.538 6.897 0.322 0.124 638.1 0.167 0.615 7.691 0.215 0.100 633.0 0.243
Ours (+ AL only) 0.549 6.566 0.291 0.105 718.0 0.183 0.629 7.880 0.176 0.076 498.9 0.206
Ours (+ AP only) 0.536 6.767 0.308 0.120 628.5 0.165 0.694 4.303 0.152 0.069 541.6 0.153
Ours (Full, AL+AP) 0.543 6.249 0.278 0.099 603.1 0.150 0.702 4.261 0.131 0.057 405.8 0.132
Improvements -0.4% 6.8% 7.3% 18.2% 12.4% 14.8% 7.7% 25.3% 32.8% 34.5% 9.3% 23.7%
Note: For a fair OCR metrics comparison, all evaluations are performed at the original input resolution. The dewarped images from DDCP are upscaled and those from LADoc are downscaled to match this resolution before metric calculation.
Figure 5: Visualization and comparison of the AD and AAD metrics on a example. (a) The ground truth image (top) and its corresponding gradient map (bottom). (b) Dewarped results from (Verhoeven et al. 2023) (top) and our method (bottom). (c) Heatmaps of the AD metric for the results shown in (b). (d) Corresponding heatmaps of our proposed AAD metric.